OpenStudy (anonymous):

plz help will give medal Use the distributive property to rewrite the expression and solve. 9*72 = 9* ______+ 9 2 =_______

OpenStudy (anonymous):

Hmm,what do u think? @surviver101

OpenStudy (anonymous):

\[9\times72=?\]

OpenStudy (anonymous):

that would be 648

OpenStudy (anonymous):

good

OpenStudy (anonymous):

\[9\times(x)+92=648\]

OpenStudy (anonymous):

let x be the unknown

OpenStudy (anonymous):

\[9x+92=648\] \[9x=648-92\]

OpenStudy (anonymous):

\[9x=556\] \[x=556\div9\]

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):

61

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

make it in fraction

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

it's \[61\frac{ 7 }{ 9 }\]

OpenStudy (anonymous):

if u want to double check ur answer..u can do this \[9\times61\frac{ 7 }{ 9 }+92=?\]

OpenStudy (anonymous):

u will get the same answer for this also \[9\times72=?\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so do u understand? @surviver101

OpenStudy (anonymous):

yes im starting to understand now

OpenStudy (anonymous):

nice :)

OpenStudy (anonymous):

so this should be right then @MARC_ 9*72 = 9* 61 7/9+ 9 2 = 648

OpenStudy (anonymous):

or is it 9*72 = 9* 61+ 9*2 = 7/9 @MARC_

OpenStudy (anonymous):

=648 is right